OpenStudy (anonymous):

A point on the rim of a .2m radius rotating wheel has a centripetal acceleration of 4m/s^2. What is the angular velocity of a point .10 m from the center of the wheel?

OpenStudy (anonymous):

good problem! Well first, what's the formula for centripetal aceleration?

OpenStudy (anonymous):

ok so Ac = v^2/r=rw^2

OpenStudy (anonymous):

yah, so you can plug in \[ 4 \ m/s^2 = (.2 \ m) \ \omega^2 \] to find omega^2

OpenStudy (anonymous):

so 4/.2=w^2 =>20 =w^2

OpenStudy (anonymous):

\[4.5rad/s^2\]

OpenStudy (anonymous):

So that's your omega, yah? The angular velocity

OpenStudy (anonymous):

sweet thanks

OpenStudy (anonymous):

Welcome ^_^ - and you know that for any pint on the wheel omeega is the same, yah? It's the magic of radians!

OpenStudy (anonymous):

ok cool ill keep that in mind.

OpenStudy (anonymous):

^_^